A business ordered 3 keyboards and a monitor from a computer company, HiTech, and sent in a payment of $675 with the order. The keyboards cost $25 each but the cost of the monitor was missing. Using words, describe what equation the business could use to find the cost of the monitor.
Please help. I neeeed the answer to this question asap, (:

Answer by ReadingBoosters(3246)

Numerical expression:
3x + y = $675
if x = $25 then y = $675 - 3x = 675 - 75 = 600
Written expression:
The business would deduct triple the cost of one $25 keyboard from their total $675 payment to determine the cost of the monitor.